Operations Report Dec 23rd

Crew 236 Operations Report 23DEC2021

SOL: 4

Name of person filing report: Pavithra Ravi

Non-nominal systems: None

Notes on non-nominal systems: N/A


Hours run: 16 hours 2 minutes

From what time last night: 16:58

To what time this morning: 09:00

List any additional daytime hours when the generator was run: None

Solar— SOC % (Before generator is run at night): Information provided by Outpost

Station Propane Reading – % Information provided by Outpost

Ethanol Free Gasoline: Not in use

Water (loft tank): 32.5 gallons

Water Meter: 1539893 units

Water (static tank): approx. 370 gallons

Static to Loft Pump used: No

Water in GreenHab: approx. 213 gallons

Water in ScienceDome: N/A

Toilet tank emptied: No

Deimos rover used: No

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Spirit rover used: No

Hours (before EVA 1): N/A

Beginning charge (Before EVA 1): N/A

Ending charge (On return from EVA 1, before recharging): N/A

Hours (before EVA 2): N/A

Beginning charge (Before EVA 2): N/A

Ending charge (On return from EVA 2, before recharging): N/A

Currently charging: Yes

Opportunity rover used: Yes

Hours: 76.6

Beginning charge: 100%

Ending charge: 75%

Currently charging: Yes

Curiosity rover used: Yes

Hours: 155.1

Beginning charge: 100%

Ending charge: 97%

Currently charging: Yes

Notes on rovers: One of the rovers was accidentally not turned off after the EVA today. This was identified about three hours after the return from the EVA and resolved with an emergency EVA.

ATV’s Used (Honda, 300, 350.1, 350.2, 350.3): None

Reason for use: N/A

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were used today: N/A

Notes on ATVs: N/A

HabCar used and why, where? No

CrewCar used and why, where? No, but we did not report mileage when leaving Grand Junction and heading to MDRS. Mileage – 198640. Tank is full. No damage besides normal wear and tear. Oil was checked — it’s about 75% full (oil level was around the 4th of 5 dots).

General notes and comments: None

Summary of internet: Nominal

Summary of suits and radios: Nominal. The new headset works properly.

Today, the fan on one of the suits slowed down considerably about two hours into the EVA, causing the crew member’s helmet to fog up. The EVA was immediately halted, and the crew attempted to return to the Hab before the fan stopped. Unfortunately, the fan completely failed just about 10 minutes after. It was later found that the suit was only plugged in partially, resulting in an incomplete charge.

Summary of Hab operations: Nominal

Summary of GreenHab operations: Nominal

Summary of ScienceDome operations: Nominal

Summary of RAM operations: Nominal

Summary of any observatory issues: In correspondence with the

astronomy support team.

Summary of health and safety issues: Nominal

Questions, concerns and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site